OPTICAL SPECTROSCOPY OF SILICON-CARBON CLUSTERS: Si_{2}C and Si_{3}C

\underline{NEIL J. REILLY}, DAMIAN L. KOKKIN, MICHAEL C. McCARTHY, and PATRICK THADDEUS, Harvard-Smithsonian Center for Astrophysics, 60 Garden St., Cambridge, MA 02138, and School of Engineering & Applied Sciences, Harvard University, 29 Oxford St., Cambridge, MA 02138.

We report the first measurement of an electronic spectrum of Si_{2}C, observed in a jet-cooled discharge through silane, acetylene and argon, in the 380 − 410 nm wavelength range. While Si_{2}C is a highly plausible astronomical molecule, searches for its rotational transitions in the laboratory and in space are impractical at present - ab initio predictions of the rotational constants of this slightly bent species have yet to performed to within the required accuracy. By analogy with SiC_{2}, the carrier of the well-known Merrill-Sanford bands, electronic spectroscopy may provide estimates of its rotational constants and structure, thereby constraining searches for its millimeter-wave transitions. Our experiments suggest that the electronic transition has a large oscillator strength and a significant fluorescence quantum yield, making it a good candidate for optical detection in space, particularly in those carbon stars where SiC_{2} is known to be abundant. As part of a more general effort to measure the electronic spectra of small silicon-carbon clusters, several examples of which have been identified in space by radio-astronomy, we present a spectrum of Si_{3}C with a much higher S/N ratio than has been previously reported, and which is now in excellent agreement with theory.